Circular No. 8732

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2006 (624) 1 F. Marchis and M. H. Wong, University of California at Berkeley

(UCB); J. Berthier, P. Descamps, D. Hestroffer, and F. Vachier, Institut de Mécanique Céleste et de Calcul des Éphémérides, Observatoire de Paris; D. Le Mignant, W. M. Keck Observatory; and I. de Pater, UCB, report the discovery of a satellite of (624) Hektor using the Keck-II telescope and its Laser Guide Star Adaptive Optics system. Observations of this L4 Trojan minor planet (V = 14.4), recorded in H- and K-band filters on July 16.582 UT over a 20-min baseline, clearly show the presence of a moonlet companion (peak S/N = 20) at 0′′.36 (p.a. 250^{o}) from the primary. The shape of the resolved primary can be approximated by an ellipse with major and minor axes 2a = 350 km and 2b = 210 km (108 and 65 milli-arcseconds), but several individual images further indicate that the primary may have a bilobated shape. Based on the integrated brightness ratio between the moonlet and the primary of ∼ 6.5, the diameter of S/2006 (624) 1 is estimated to be ∼ 15 km.

SUPERNOVAE 2006ds AND 2006dt Two additional supernovae have been discovered on unfiltered CCD

images: 2006ds by Y.-S. Tsai and K.-C. Lu (0.40-m reflector, Kenting Star Village Observatory, Hengchun, Taiwan; note spelling correction to Lu’s name from the initial report on CBET 580), and 2006dt by N. Lee and W. Li (via LOSS/KAIT; cf. IAUC 8731).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ds July 18.80 22^{h}55^{m}18.^{s}32 −33^{o}55′17′′.5 16.1 5′′.4 W, 3′′.6 S
2006dt July 20.20 14 35 43.11 + 8 17 55.8 16.8 2′′.9 E, 6′′.2 S

SN 2006ds in PGC 70011 was confirmed at mag ∼ 15.5 on a CCD frame taken by Tsai and Lu on July 19.82 UT; nothing is visible at this location on a Digitized Sky Survey image (limiting red mag 20.5). SN 2006ds is a type-II supernova (spectroscopic details on CBET 580). Additional magnitudes for 2006dt in NGC 5681: July 11.20, [19.0; 15.23, 17.5. SN 2006dt is a type-Ia supernova, near maximum on July 21.22 (details on CBET 582).

2006 July 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green